OpenStudy (anonymous):

(x+11)2/16 - y2/9 =1 properties of the conic? center, vertices, foci ?

OpenStudy (anonymous):

\[\frac{ (x+11)^{2} }{ 16 }-\frac{ y^{2} }{ 9 } =1\] find the properties.

OpenStudy (anonymous):

X^2+22X+121/16 - Y^2/9 =